OpenStudy (anonymous):

How many milliliters of 0.149 M HCl would be required to titrate 5.58 g KOH?

OpenStudy (gebooors):

Neutralization reaction: HCl (aq) + KOH (aq) -> H2O (l) + KCl (aq) Equal amounts of HCl and KOH neutralizes each other. First calculate n (KOH) = m/M = n(HCl) Then solve from equation n =cV volume V 0,149 M = 0,149 mol/l You get volume in litres.